At the supermarket Cammy can buy 2/5 pound of oranges for $4. At the farmer's market she can buy 4/7 pound of oranges for the same price. Where does she get the better deal?
Answer by Boreal(15235)

2/5 of a pound for $4
multiply both by 5/2 Then 1 pound is $20/2=$10.
-----------------
4/7 pound for $4
multiply both by 7/4=$28/4=$7 per pound.
The farmer's market.